Energy Balance

\underline{Problem 1}: Given a saturated liquid stream of 30 gmol/s propane and 70 gmol/s 1-butene at 10 bar • Find the vapor fraction of this stream if it is throttled down to 2 bar.

• Find the heat load to vaporize 60% of the stream at 10 bar. Solve the above problem using a spread-sheet as well as in CHEMCAD.